OpenStudy (anonymous):

So solving equations backwards is not my thing...Write a word problem that can be solved by working backward. Include at least two steps in the problem. Then, show the answer and how the problem can be solved by working backward.

OpenStudy (anonymous):

Give it a try, we will help. Make a name for your person and state that he used half of his money to get into Disneyland and then bought ...

OpenStudy (anonymous):

Any luck?

OpenStudy (anonymous):

i just got it to pull it up lol:)

OpenStudy (anonymous):

no worries ...

OpenStudy (anonymous):

and yes it gave my an idea:)

OpenStudy (anonymous):

we can start with the problem you asked earlier, now that it is clarified Emily brought some money on a trip to the aquarium. She used half of the money to buy an admission ticket. Then, she $8 used to buy lunch at the food court. She used 3/4 of what was left to buy a souvenir, leaving her with $4 . How much money did she start with?

OpenStudy (anonymous):

lets work it out step by step, starting by defining a variable x = amount of money she came to the aquarium with.

OpenStudy (anonymous):

then the statement She used half of the money to buy an admission ticket. means she spent half of her original amount, or \[0.5x\] leaving her with \[x-0.5x=0.5x\]

OpenStudy (anonymous):

Then, she $8 used to buy lunch at the food court. we can translate that into \[.5x-8\]

OpenStudy (anonymous):

She used 3/4 of what was left to buy a souvenir, leaving her with 1/4 of what remained, or \[\frac{1}{4}(.5x-8)=\frac{1}{8}x-2\]

OpenStudy (anonymous):

leaving her with $4 \[\frac{1}{8}x-2=4\] and that is the equation we need to solve

OpenStudy (anonymous):

takes two steps only, 1) add 2 to both sides 2) multiply by 8 we get \[\frac{1}{8}x-2=4\] \[\frac{1}{8}x=6\] \[x=6\times 8=48\]

OpenStudy (anonymous):

now you can mimic this to get your equation, perhaps just change some numbers.